a rectangular prism has height 4 millimeters and width 5 millimeters.If the total surface area is 166 square millimeters, what is the length of the prism
Answer by bucky(2189)

The rectangular prism has two ends that are each 4 mm by 5 mm. The product of 4 mm times 5 mm
is 20 sq mm and since there are two of these ends, the total surface area of the ends is 40 sq mm.
.
This means that the remaining surface area is 166 sq mm minus 40 sq mm = 126 sq mm and this is
the surface area that is left for the sides. If L represents the length of the prism, then
there are two sides that each have a surface area of 4 mm times L and two sides that each have
a surface area of 5 mm times L. That means that the remaining surface area can be calculated from:
.
2(4L + 5L) = 2(9L) = 18L
.
So:
.
18L = 126 sq mm
.
Solve for L by dividing both sides by 18 and you get:
.
L = 126/18 = 7 mm
.
Hope this shows you a way through the problem. It probably would be a little more understandable
if you made a sketch.
.